OpenStudy (anonymous):

Eliminate the parameter x=t-3, y=t^2+5

OpenStudy (rogue):

x=t-3, y=t^2+5 t = x + 3 ---> y = t^2 + 5 y = (x + 3)^2 + 5 y = x^2 + 6x + 14